# Long-range Entanglement and Role of Realistic Interaction in Braiding of Non-Abelian Quasiholes in Fractional Quantum Hall Phases

### v2

@@ IV.2 Effect of quasihole energetics on braiding
We stress that while the effect of ground-state degeneracy on statistics of MR quasiholes have been discussed before, past literature focused on the effect of fusion channel selection of the same pair of $\sigma$-quasiholes being braided[44, 45]. The fusion in that context can then be thought of as a “virtual” process which helps to determine the braiding phase via the spin-statistics relation for anyons[56, 57]. Here, we show that the fusion of one pair of quasiholes in real space (and real time) can affect the braiding of other quasiholes in the system. This phenomenon may present many challenges in experiments, as the configuration of background quasiholes may be heavily influenced by impurities in the sample, which is beyond control.
@@ IV.3 Tuning the non-Abelian gap in disordered systems
We now turn attention to the “non-Abelian gap”, which we define as the energy gap between different orthogonal states within the non-Abelian degeneracy with respect to some two-body electron interaction. Non-Abelian statistics can only be realized when this gap is much smaller than the thermal energy (which in turn cannot be so large that topological phase is completely destroyed). This motivates our discussion on tuning the electron interaction to realize this limit. To simplify the discussion, we consider in this section the system with four MR quasiholes, where a two-fold degeneracy is expected when the positions of all quasiholes are fixed by the one-body potential. However, if a charge-$e/2$ is present anywhere in the system, then this degeneracy is split by the effect of two-body interaction as discussed in the section above. The toy-model Hamiltonian used in this section is
@@
Figure 7: (a) Energy gap between the lowest two eigenstates of the Hamiltonian given by Eq.(29) plotted against $\lambda$ for $v_{0}=0.1$ (blue), $v_{0}=1.0$ (orange), and $v_{0}=2.0$. The total strength of two-body interactions is normalized to unity, so $v_{0}$ represents relative strength of the one-body potential compared to two-body interaction. Numerical data were otained by exact diagonalization in the $L_{z}=0$ sector for a MR system consisting of 12 electrons and 24 orbitals (b)-(d) The overlaps of the ground state of Eq.(29) with the model state of two $1$-type anyons (solid lines) and with the model state of two $\psi$-type anyons (dashed lines), and the total overlap with these two states (dot-dashed lines) plotted against $\lambda$ for (b) $v_{0}=0.1$ (c) $v_{0}=0.5$ (d) $v_{0}=2.0$. Boxed overlays denotes a parameter region ideal for detecting non-Abelian statistics.
While this interaction is purely artificial, our analysis serves as a proof-of-concept that in principal it is possible to make the non-Abelian gap approach zero by tuning the two-body interaction. Our results also suggest a range for the ratio between the amounts of $\hat{V}_{1}^{2bdy}$ and $\hat{V}_{3}^{2bdy}$ in order to achieve this. This can serve as a rough estimation for the target range of electron-electron interaction for stabilizing non-Abelian properties in experiments. In the supplementary material, we also relate this toy model to realistic interactions, namely the screened Coulomb interaction and the Zhang-Das Sarma (ZDS) interaction modelling the effect of finite sample thickness[80]. Preliminary results of numerical simulations will be presented regarding the possibility of using screening strength and sample thickness as tuning parameters in experiment to stabilize the non-Abelian degeneracy.
@@
$\hat{V}_{3}^{2bdy}$ Longer-ranged (“hollow-core”) electron-electron interaction prefers $1$-type quasihole
@@
We can see the effect of the interplay between these three potentials in Fig. 7a, which shows the gap between the lowest two eigen-energies of Eq. (29) for a system consisting of 12 electrons and 24 orbitals. With this system there are four MR quasiholes, which pair-wise form one charge $e/2$ at the north pole and another at the south pole. In this choice of potential pins configuration the system exhibit rotational symmetry about the $z$-axis, which allows us to further reduce the the necessary Hilbert space used in numerics from the full MR CHS to only the $L_{z}=0$ sector.
@@
In Fig. 7a one can observe the combined effect of everything we have learned about quasihole energetics. If the potential pin strength $v_{0}$ is too large, such that it dominates over the two-body interaction, then the ground state is always the “$1+1$” state (given by Eq.(15)), which consists of two $1$-anyons. However, if the trapping pins are not so trong and the two-body interaction is dominated by $\hat{V}_{1}^{2bdy}$, it is possible that the ground state is “$\psi+\psi$” (given by Eq.(16)) instead. In this weak pinning potential limit, if one starts adding in $\hat{V}_{3}^{2bdy}$ by tuning $\lambda$ in Eq. (29), the Hamiltonian will start to favor the “$1+1$” state again. In the limit $\lambda\to 1$ the Hamiltonian completely favors the “$1+1$” state agains since both $\hat{V}_{3}^{2bdy}$ and $\hat{V}_{\text{pins}}$ favor it. Thus, there must be a point in the middle of the two limits where a crossing between the two states happen. We can see this crossing point in Fig. 7b-c (it is also non-existent in Fig. 7d in the strong potential pin limit). It is noteworthy that the gap does not always fully close at the crossing point, which could be a band repulsion effect coming from anyon-anyon interaction in a finite system. However, at the crossing point the gap reaches its minimum, which can be either vanishing in the thermodynamic limit, or potentially overcome by thermal energy in real experiments. Most notably the total overlap with the two model states (defined by the root sum squared of the two overlaps) is very close to unity near the crossing point. This indicates that the two lowest, near-degenerate states are fully described by the model states for the $1$-type and $\psi$-type anyon. Potentially, this means that this parameter range (denoted by an boxed overlay in Fig. 7c) is ideal for carrying out non-Abelian braiding. However, to be able to confidently make this statement requires further analysis into scenarios also involving charge-$e/4$ anyons, which will be presented elsewhere.
@@ Acknowledgement
S1 More on parity conservation
S1.1 Jack polynomials for the MR states
@@
This binary string can be treated as a representation of electrons occupying the Landau orbital. Each digit corresponds to an orbital, which on the spherical geometry used throughout this paper corresponds to angular momentum quantum numbers $L_{z}=S,S-1,S-2,...,-S+2,-S+1,-S$ where $S$ is an integer or half-integer. A digit $1$ means that the orbital is occupied by an electron while a digit 00 means it is empty. As per our convention, the left-most orbital corresponds to the north pole ($L_{z}=S$) and the right-most orbital corresponds to the south pole ($L_{z}=-S$).
One can check that Eq.(S1) satisfies the $(2,4)$-admissibility rule: there exists no more than two electrons within four consecutive orbitals (i.e. no more than two $1$’s within any four conseuctive digits). Any $(2,4)$-admissible configuration can serve as a “root configuration” for a Jack polynomial that normalizes to the MR state. Given a root configuration and an accompanying rational parameter $\alpha$, which equals to $-2$ for MR states, the monomial coefficients of a Jack polynomial can be computed recursively[65, 66, 81].
Let us denote $N_{o}$ the number of orbitals (number of digits) and $N_{e}$ the number of electrons (number of $1$’s). We can see that Eq.(S1) satisfies the commensurability condition $N_{o}=2N_{e}-2$. This is the unique root configuration that minimizes the ratio $N_{e}/N_{o}$ while satisfying the $(2,4)$-admissible rule — there exists no $(2,4)$-admissible ocnfiguration with $N_{o}<2N_{e}-2$. Jack polynomials with such root configurations give the ground state of the MR state. Quasihole states are obtained by Jacks whose root configurations have $N_{o}>2N_{e}-2$. If a positive integer $n$ is the number of added orbitals to the ground state, i.e. $N_{o}=2N_{e}-2+n$, then $2n$ quasiholes are created.
S1.2 Parity conservation seen from circles and crosses
@@
where the ellipses “...” denotes repeating patterns of $1100$. Afterward, all other admissible root can be generated recursively by picking a digit $1$ and move it one unit to the left, provided both that the digit to its left is 00 and that the resulting configuration does not violate the admissible rule. Following this procedure, the first few terms go like
@@
[…]
A quasihole is marked at the center of each group of four digits with either one symbol (if the group contains one $1$) or two symbols (if the group contains two $1$’s). The choice of symbol can either be circles ($\circ$”) or crosses (“$\times$”) based on the following rule: we first divide $2n$ quasiholes into $n$ pairs, going from left to write, then mark each pair with circles if there are an even number of $1$’s between them, or crosses if there are an odd number of $1$’s between them.
@@
[…]
@@
  • In the rule for marking circles and crosses, we divide the quasiholes into pairs based on their order of appearance from left to right. A priori, this is a rather arbitrary choice as in general a fusion channel exists for any pair of quasiholes. However, one can further check that no matter how $2n$ quasiholes are divided into $n$ pairs, as long as we mark each pair as two circles and two crosses based on whether the number of $1$’s between them is even or odd, respectively, the resulting numbers of circles and crosses are the same. Thus, the exact marking scheme does not affect our parity conservation, which only counts the total number of crosses regardless of where they occur in the root configuration. In other words, parity is well-defined in our “circles-and-crosses” description.
@@
S2 Fusion channel selection and braiding statistics from tree diagrams
S2.1 Fusion Tree Diagram - A very brief introduction
@@
The fusion rule $\sigma\times\sigma=1+\psi$ means two diagrams are possible, corresponding to the even sector (left) and odd sector (right) respectively. Each diagram are to be read from bottom to top. Each diagram in Eq.(S21) represents the process of creating a pair of $\sigma$-anyons from vacuum (“$1$”) or from a single Majorana fermion (“$\psi$”). A four-$\sigma$ state can be similarly represented, respectively in the even and odd sector, by the following diagrams:
@@
Each diagram in Eq.(S22) represents the creation of four $\sigma$ quasiholes from three pair creation processes (also called “splitting”). In the second splitting, the fusion rule allows for two possible outcomes, so each diagram actually represents a two-fold degenerate space of four quasiholes. A general diagram representing $2n$ $\sigma$-anyons is as follows
@@
for the even sector (for the odd sector, the bottom $1$ is replaced with $\psi$). The degeneracy of this space can also be read off from the diagram: $2n$ quasiholes are created by $12n-$ 1 processes, $n-1$ of which independently have two possible outcomes, thus the total degeneracy is $2^{n-1}$.
S2.2 Effect of fusion channel selection on anyon statistics
@@
Here, after the fusion of $\sigma_{5}$ and $\sigma_{6}$, the ground state manifold consisting of $\sigma_{1}$, $\sigma_{2}$, $\sigma_{3}$, $\sigma_{4}$, and the fused quasihole is still two-fold degenerate. Thus, the four $\sigma$-anyons still exhibit non-Abelian statistics regardless of the fusion outcome preferred by the dynamics. However, due to the different “sub-diagrams”, the two fusion outcomes will results in different Berry matrices[47].
S3 Creation energy of MR quasiholes
The self-energy of $\sigma$-anyons can never be determined directly because of the simultaneous creation of two $\sigma$-anyons from one flux insertion in MR non-Abelian phases. But it can be asymptotically computed by the method we shall explain here. The expectation value of any electron-electron interaction regarding the state with two $\sigma$-anyons has three components: the ground-state energy, two quasihole creation energy (effective self-energy), and the interaction between the two quasiholes. Considering the state represented by Eq. (8) where two $\sigma$-anyons are positioned at the North and South pole, the separation between them ($2R$ as $R$ being the radius of the sphere) is enlarged when we increase the system’s size. When $R$ (namely the separation between two anyons) is large enough, there should be no interaction between them and only self-energies remain. The energies of the state in Eq. (8) with the model Hamiltonian $\hat{V}_{1}$, $\hat{V}_{3}$, and $\hat{V}_{5}$ subtracted by the ground-state energies are shown in Figure S1 for both odd and even numbers of electrons. Though the anyon types are both $\sigma$, their effective self-energies in the two different sectors are different at least up to the finite system sizes we could obtain ($18$ electrons for $1$-sector and $19$ electrons for $\psi$-sector).
Figure S1: Energy difference between the NN state and the NS state with different pseudopotentials: (a) with $\hat{V}_{1}$, (b) with $\hat{V}_{3}$, (c) with $\hat{V}_{5}$. The systems’ sizes are from 10 electrons to 18 electrons, increased every time by two for even and from 11 electrons to 19 electrons increased every time by two for odd.
S4 Layer Thickness and Screening as Tuning Parameters
@@
S4.1 Finite thickness effect
@@
S5 Details on the numerical methods
@@
S5.1 General Hamiltonian and numerical constrains
@@
where $\hat{V}_{m}^{nbdy}$ denotes the $m$-th $n$-body pseudopotential, $\lambda_{0}$ and $\lambda_{1}$ are some scalars, and $\hat{V}_{\text{pins}}$ is a one-body pinning potential that depends on some fictituous “time” parameter $t$. The exact form of $\hat{V}_{\text{pins}}$ will be described later.
@@
The system consists of two potential pins positioned opposite of each other on a small circle parallel to the equator (first two terms) and a third pin at the south pole (third term) (see inset of Fig. 7a in the main text). The pin at the south pole is twice the strength of each of the other two pins, so that when $\theta_{0}=0$ it is equivalent to two identical pin at the north and south poles. The polar angle $\phi_{0}(t)$ is varied between $[0,\pi]$ from $t=0$ to $t=T$ so that the two pins are exchanged at the end of the process.
Figure S3: The spectrum of Eq.(S30) plotted against parameter $\theta_{0}$ for MR systems with 10 electrons (top row) and 11 electrons (bottom row) with $\lambda_{1}=0.1$ (first column), $\lambda_{1}=1.0$ (second column), and $\lambda_{1}=5.0$ (third column).
@@
In our numerics, we empirically choose $\lambda_{1}=5.0$, which is large so that the spectrum is qualitatively different (see Fig. S3) but not so large that the quasihole interaction dominates. However, it is interesting to note in Fig. S3 that for all values of $\lambda_{1}$ shown, there exist more than two states with low-lying energy. The two lowest energy states are the states of interest—consisting of two charge-$e/4$ anyons and one charge-$e/2$ anyons of either $1$-type or $\psi$-type; the remaining are various other anyon configurations. Especially at finite temperature, all of these states may be important for understanding experimental outcomes. However, their nature remains to be understood.
S5.2 Berry phase and braiding phase
@@
where the $t_{i}$’s are discretized time: $0=t_{0}<t_{1}<t_{2}<...<t_{N}=T$. In the limit $N\to\infty$, $\mathcal{D}\to 1$ and $\gamma\to\gamma_{\text{Berry}}$ is the desired Berry phase. In practice, $N$ is chosen large enough such that $\mathcal{D}$ is close to unity, and the closeness of $\mathcal{D}$ to unity indicates how accurate the determined value of $\gamma$ is to the real value $\gamma_{\text{Berry}}$. Eq. (S37) also has an advantage of being gauge-independent. This is specially useful for numerics where $|\psi(t_{i})\rangle$ is obtained by exact diagonalization (ED), since any ED algorithm can only determine the eigenvector up to a random phase factor. However, we can see in Eq. (S37) each ket has a corresponding bra, thus cancelling out the effect of such random phases.

### v3

@@ IV.2 Effect of quasihole energetics on braiding
We stress that while the effect of ground-state degeneracy on statistics of MR quasiholes have been discussed before, past literature focused on the effect of fusion channel selection of the same pair of $\sigma$-quasiholes being braided[44, 45]. The fusion in that context can then be thought of as a ``virtual'' process which helps to determine the braiding phase via the spin-statistics relation for anyons[56, 57]. Here, we show that the fusion of one pair of quasiholes in real space (and real time) can affect the braiding of other quasiholes in the system. This phenomenon may present many challenges in experiments, as the configuration of background quasiholes may be heavily influenced by impurities in the sample, which is beyond control.
@@ IV.3 Tuning the non-Abelian gap in disordered systems
We now turn attention to the ``non-Abelian gap'', which we define as the energy gap between different orthogonal states within the non-Abelian degeneracy with respect to some two-body electron interaction. Non-Abelian statistics can only be realized when this gap is much smaller than the thermal energy (which in turn cannot be so large that topological phase is completely destroyed). This motivates our discussion on tuning the electron interaction to realize this limit. To simplify the discussion, we consider in this section the system with four MR quasiholes, where a two-fold degeneracy is expected when the positions of all quasiholes are fixed by the one-body potential. However, if a charge-$e/2$ is present anywhere in the system, then this degeneracy is split by the effect of two-body interaction as discussed in the section above. The toy-model Hamiltonian used in this section is
@@
Figure 8: (a) Energy gap between the lowest two eigenstates of the Hamiltonian given by Eq.(35) plotted against $\lambda$ for $v_{0}=0.1$ (blue), $v_{0}=1.0$ (orange), and $v_{0}=2.0$. The total strength of two-body interactions is normalized to unity, so $v_{0}$ represents relative strength of the one-body potential compared to two-body interaction. Numerical data were otained by exact diagonalization in the $L_{z}=0$ sector for a MR system consisting of 12 electrons and 24 orbitals (b)-(d) The overlaps of the ground state of Eq.(35) with the model state of two $1$-type anyons (solid lines) and with the model state of two $\psi$-type anyons (dashed lines), and the total overlap with these two states (dot-dashed lines) plotted against $\lambda$ for (b) $v_{0}=0.1$ (c) $v_{0}=0.5$ (d) $v_{0}=2.0$. Boxed overlays denotes a parameter region ideal for detecting non-Abelian statistics[79].
@@
$\hat{V}_{3}^{2bdy}$ Longer-ranged (``hollow-core'') electron-electron interaction prefers $1$-type quasihole
@@
We can see the effect of the interplay between these three potentials in Fig. 8a, which shows the gap between the lowest two eigen-energies of Eq. (35) for a system consisting of 12 electrons and 24 orbitals (similar trends are obtained for systems with 8 and 10 electrons, as shown in the supplementary materials). With this system there are four MR quasiholes, which pair-wise form one charge $e/2$ at the north pole and another at the south pole. In this choice of potential pins configuration the system exhibit rotational symmetry about the $z$-axis, which allows us to further reduce the the necessary Hilbert space used in numerics from the full MR CHS to only the $L_{z}=0$ sector.
@@
In Fig. 8a one can observe the combined effect of everything we have learned about quasihole energetics. If the potential pin strength $v_{0}$ is too large, such that it dominates over the two-body interaction, then the ground state is always the ``$1+1$'' state (given by Eq.(15)), which consists of two $1$-anyons. However, if the trapping pins are not so trong and the two-body interaction is dominated by $\hat{V}_{1}^{2bdy}$, it is possible that the ground state is $\psi+\psi''$ (given by Eq.(16)) instead. In this weak pinning potential limit, if one starts adding in $\hat{V}_{3}^{2bdy}$ by tuning $\lambda$ in Eq. (35), the Hamiltonian will start to favor the ``$1+1$'' state again. In the limit $\lambda\to 1$ the Hamiltonian completely favors the ``$1+1$'' state agains since both $\hat{V}_{3}^{2bdy}$ and $\hat{V}_{\text{pins}}$ favor it. Thus, there must be a point in the middle of the two limits where a crossing between the two states happen. We can see this crossing point in Fig. 8b-c (it is also non-existent in Fig. 8d in the strong potential pin limit). It is noteworthy that the gap does not always fully close at the crossing point, which could be a band repulsion effect coming from anyon-anyon interaction in a finite system. However, at the crossing point the gap reaches its minimum, which can be either vanishing in the thermodynamic limit, or potentially overcome by thermal energy in real experiments. Most notably the total overlap with the two model states (defined by the root sum squared of the two overlaps) is very close to unity near the crossing point. This indicates that the two lowest, near-degenerate states are fully described by the model states for the $1$-type and $\psi$-type anyon. Potentially, this means that this parameter range (denoted by an boxed overlay in Fig. 8c) is ideal for carrying out non-Abelian braiding. However, to be able to confidently make this statement requires further analysis into scenarios also involving charge-$e/4$ anyons, which will be presented elsewhere.
While the interaction in Eq. (35) is purely artificial, our analysis serves as a proof-of-concept that in principal it is possible to make the non-Abelian gap approach zero by tuning the two-body interaction. Our results also suggest a range for the ratio between the amounts of $\hat{V}_{1}^{2bdy}$ and $\hat{V}_{3}^{2bdy}$ in order to achieve this. This can serve as a rough estimation for the target range of electron-electron interaction for stabilizing non-Abelian properties in experiments. In real experiments, this ratio can be tuned by tuning the range of the effective electron-electron interactions. Roughly speaking, a longer-ranged interaction implies higher $\hat{V}_{3}^{2bdy}$-to-$\hat{V}_{1}^{2bdy}$ ratio. This is possible by one of the two methods: either tuning the finite sample thickness, or tuning the strength of the screening of the Coulomb interaction. In the supplementary material, we provide a brief quantitative discussion on the effects of these two parameters, using the the Zhang-Das Sarma (ZDS) interaction to model the effect of finite sample thickness[83] and the Yukawa potential to model Coulomb screening.
@@ Acknowledgement
S6 More on parity conservation
S6.1 Jack polynomials for the MR states
@@
This binary string can be treated as a representation of electrons occupying the Landau orbital. Each digit corresponds to an orbital, which on the spherical geometry used throughout this paper corresponds to angular momentum quantum numbers $L_{z}=S,S-1,S-2,...,-S+2,-S+1,-S$ where $S$ is an integer or half-integer. A digit $1$ means that the orbital is occupied by an electron while a digit 0 means it is empty. As per our convention, the left-most orbital corresponds to the north pole ($L_{z}=S$) and the right-most orbital corresponds to the south pole ($L_{z}=-S$).
One can check that Eq.(S36) satisfies the $(2,4)$-admissibility rule: there exists no more than two electrons within four consecutive orbitals (i.e. no more than two $1$'s within any four conseuctive digits). Any $(2,4)$-admissible configuration can serve as a ``root configuration'' for a Jack polynomial that normalizes to the MR state. Given a root configuration and an accompanying rational parameter $\alpha$, which equals to 2-2 for MR states, the monomial coefficients of a Jack polynomial can be computed recursively[65, 66].
Let us denote $N_{o}$ the number of orbitals (number of digits) and $N_{e}$ the number of electrons (number of $1$'s). We can see that Eq.(S36) satisfies the commensurability condition $N_{o}=2N_{e}-2$. This is the unique root configuration that minimizes the ratio $N_{e}/N_{o}$ while satisfying the $(2,4)$-admissible rule — there exists no $(2,4)$-admissible ocnfiguration with $N_{o}<2N_{e}-2$. Jack polynomials with such root configurations give the ground state of the MR state. Quasihole states are obtained by Jacks whose root configurations have $N_{o}>2N_{e}-2$. If a positive integer $n$ is the number of added orbitals to the ground state, i.e. $N_{o}=2N_{e}-2+n$, then $2n$ quasiholes are created.
S6.2 Parity conservation seen from circles and crosses
@@
where the ellipses ``...'' denotes repeating patterns of $1100$. Afterward, all other admissible root can be generated recursively by picking a digit $1$ and move it one unit to the left, provided both that the digit to its left is 0 and that the resulting configuration does not violate the admissible rule. Following this procedure, the first few terms go like
@@
 11   $1\displaystyle\hbox{\hskip 2.5pt\hskip-2.5pt\hbox{$\color[rgb]{1,1,1}{1}$}\hskip-2.5pt\hskip-5.7083pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu\scriptstyle\circ{\color[rgb]{1,1,1}{1}}$}}\hskip-5.7083pt\hskip 2.5pt}\hbox{\hskip 2.5pt\hskip-2.5pt\hbox{$\hbox{\hskip 2.5pt\hskip-2.5pt\hbox{$0$}\hskip-2.5pt\hskip-4.65277pt\raisebox{-9.55556pt}{\hbox{$\mkern 2.0mu\scriptstyle{\circ\circ}$}}\hskip-4.65277pt\hskip 2.5pt}$}\hskip-2.5pt\hskip-4.65277pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu\scriptstyle{\circ\circ}$}}\hskip-4.65277pt\hskip 2.5pt}0\hbox{\hskip 2.5pt\hskip-2.5pt\hbox{$0$}\hskip-2.5pt\hskip-5.7083pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu\scriptstyle\circ{\color[rgb]{1,1,1}{1}}$}}\hskip-5.7083pt\hskip 2.5pt}1100110011...11001$1 (S50)
@@
A quasihole is marked at the center of each group of four digits with either one symbol (if the group contains one $1$) or two symbols (if the group contains two $1$'s). The choice of symbol can either be circles ($\circ''$) or crosses ($\times''$) based on the following rule: we first divide $2n$ quasiholes into $n$ pairs, going from left to write, then mark each pair with circles if there are an even number of $1$'s between them, or crosses if there are an odd number of $1$'s between them.
@@
     1 $\hbox{\hskip 2.5pt\hskip-2.5pt\hbox{$\color[rgb]{1,1,1}{1}$}\hskip-2.5pt\hskip-4.65277pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu\scriptstyle{\circ\circ}$}}\hskip-4.65277pt\hskip 2.5pt}01100110011...110011$ (S54)
     1×× $\hbox{\hskip 2.5pt\hskip-2.5pt\hbox{$1$}\hskip-2.5pt\hskip-5.97227pt\raisebox{-4.7222pt}{\hbox{$\mkern 2.0mu\scriptscriptstyle{\times\times}$}}\hskip-5.97227pt\hskip 2.5pt}001100110011...11001$1 (S55)
@@
  • In the rule for marking circles and crosses, we divide the quasiholes into pairs based on their order of appearance from left to right. A priori, this is a rather arbitrary choice as in general a fusion channel exists for any pair of quasiholes. However, one can further check that no matter how $2n$ quasiholes are divided into $n$ pairs, as long as we mark each pair as two circles and two crosses based on whether the number of $1$'s between them is even or odd, respectively, the resulting numbers of circles and crosses are the same. Thus, the exact marking scheme does not affect our parity conservation, which only counts the total number of crosses regardless of where they occur in the root configuration. In other words, parity is well-defined in our ``circles-and-crosses'' description.
@@
S7 Fusion channel selection and braiding statistics from tree diagrams
S7.1 Fusion Tree Diagram - A very brief introduction
@@
The fusion rule $\sigma\times\sigma=1+\psi$ means two diagrams are possible, corresponding to the even sector (left) and odd sector (right) respectively. Each diagram are to be read from bottom to top. Each diagram in Eq.(S56) represents the process of creating a pair of $\sigma$-anyons from vacuum (``$1$'') or from a single Majorana fermion ($\psi''$). A four-$\sigma$ state can be similarly represented, respectively in the even and odd sector, by the following diagrams:
@@
Each diagram in Eq.(S57) represents the creation of four $\sigma$ quasiholes from three pair creation processes (also called ``splitting''). In the second splitting, the fusion rule allows for two possible outcomes, so each diagram actually represents a two-fold degenerate space of four quasiholes. A general diagram representing $2n$ $\sigma$-anyons is as follows
@@
for the even sector (for the odd sector, the bottom $1$ is replaced with $\psi$). The degeneracy of this space can also be read off from the diagram: $2n$ quasiholes are created by 2n12n-1 processes, n1n-1 of which independently have two possible outcomes, thus the total degeneracy is $2^{n-1}$.
S7.2 Effect of fusion channel selection on anyon statistics
@@
Here, after the fusion of $\sigma_{5}$ and $\sigma_{6}$, the ground state manifold consisting of $\sigma_{1}$, $\sigma_{2}$, $\sigma_{3}$, $\sigma_{4}$, and the fused quasihole is still two-fold degenerate. Thus, the four $\sigma$-anyons still exhibit non-Abelian statistics regardless of the fusion outcome preferred by the dynamics. However, due to the different ``sub-diagrams'', the two fusion outcomes will results in different Berry matrices[47].
S8 Creation energy of MR quasiholes
The self-energy of $\sigma$-anyons can never be determined directly because of the simultaneous creation of two $\sigma$-anyons from one flux insertion in MR non-Abelian phases. But it can be asymptotically computed by the method we shall explain here. The expectation value of any electron-electron interaction regarding the state with two $\sigma$-anyons has three components: the ground-state energy, two quasihole creation energy (effective self-energy), and the interaction between the two quasiholes. Considering the state represented by Eq. (8) where two $\sigma$-anyons are positioned at the North and South pole, the separation between them ($2R$ as $R$ being the radius of the sphere) is enlarged when we increase the system's size. When $R$ (namely the separation between two anyons) is large enough, there should be no interaction between them and only self-energies remain. The energies of the state in Eq. (8) with the model Hamiltonian $\hat{V}_{1}$, $\hat{V}_{3}$, and $\hat{V}_{5}$ subtracted by the ground-state energies are shown in Figure S1 for both odd and even numbers of electrons. Though the anyon types are both $\sigma$, their effective self-energies in the two different sectors are different at least up to the finite system sizes we could obtain ($18$ electrons for $1$-sector and $19$ electrons for $\psi$-sector).
Figure S1: Energy difference between the $N$ state and the NS state with different pseudopotentials: (a) with $\hat{V}_{1}$, (b) with $\hat{V}_{3}$, (c) with $\hat{V}_{5}$. The systems' sizes are from 10 electrons to 18 electrons, increased every time by two for even and from 11 electrons to 19 electrons increased every time by two for odd.
S9 Layer Thickness and Screening as Tuning Parameters
@@
S9.1 Finite thickness effect
@@
S10 Details on the numerical methods
@@
S10.1 General Hamiltonian and numerical constrains
@@
where $\hat{V}_{m}^{nbdy}$ denotes the $m$-th $n$-body pseudopotential, $\lambda_{0}$ and $\lambda_{1}$ are some scalars, and $\hat{V}_{\text{pins}}$ is a one-body pinning potential that depends on some fictituous ``time'' parameter $t$. The exact form of $\hat{V}_{\text{pins}}$ will be described later.
@@
The system consists of two potential pins positioned opposite of each other on a small circle parallel to the equator (first two terms) and a third pin at the south pole (third term) (see inset of Fig. 7a in the main text). The pin at the south pole is twice the strength of each of the other two pins, so that when $\theta_{0}=0$ it is equivalent to two identical pin at the north and south poles. The polar angle $\phi_{0}(t)$ is varied between $0,\pi]$ from $t=0$ to $t=T$ so that the two pins are exchanged at the end of the process.
@@
In our numerics, we empirically choose $\lambda_{1}=5.0$, which is large so that the spectrum is qualitatively different (see Fig. 6 in the main text) but not so large that the quasihole interaction dominates. As pointed out in the main text, there exist more than two states with low-lying energy. The two lowest energy states are the states of interest—consisting of two charge-$e/4$ anyons and one charge-$e/2$ anyons of either $1$-type or $\psi$-type; the remaining are various other anyon configurations. Especially at finite temperature, all of these states may be important for understanding experimental outcomes. However, their nature remains to be understood.
S10.2 Berry phase and braiding phase
@@
where the $t_{i}'$s are discretized time: $0=t_{0}<t_{1}<t_{2}<...<t_{N}=T$. In the limit $N\to\infty$, $\mathcal{D}\to 1$ and $\gamma\to\gamma_{\text{Berry}}$ is the desired Berry phase. In practice, $N$ is chosen large enough such that $\mathcal{D}$ is close to unity, and the closeness of $\mathcal{D}$ to unity indicates how accurate the determined value of $\gamma$ is to the real value $\gamma_{\text{Berry}}$. Eq. (S72) also has an advantage of being gauge-independent. This is specially useful for numerics where $|\psi(t_{i})\rangle$ is obtained by exact diagonalization (ED), since any ED algorithm can only determine the eigenvector up to a random phase factor. However, we can see in Eq. (S72) each ket has a corresponding bra, thus cancelling out the effect of such random phases.